OpenStudy (anonymous):

find all solutions for cos z = 2, z=x+iy?

OpenStudy (anonymous):

2-iy = x

OpenStudy (anonymous):

2-x/i = y

OpenStudy (anonymous):

2-x/y = i

OpenStudy (anonymous):

you need to start with \[\frac{1}{2}(e^{iz}+e^{-iz})=2\]

OpenStudy (anonymous):

then to \[e^{iz}+e^{-iz}=4\] and then \[e^{2iz}+4e^{iz}+1=0\] solve the quadratic equation in \(e^{iz}\)

OpenStudy (anonymous):

typo i mean \[e^{2iz}-4e^{iz}+1=0\]

OpenStudy (anonymous):

ohh. ic. thx.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

cos never be 2 in maths

OpenStudy (anonymous):

-1<cosx<1